# Molecular Fingerprints Are Strong Models for Peptide Function Prediction

### v1

@@ 4.1 Long Range Graph Benchmark (LRGB)
To begin, we report the results on the LRGB (Dwivedi et al., 2022) benchmark, on both Peptides-func (classification) and Peptides-struct (regression), in Table 1. Those are multioutput datasets, and we report the metric values averaged from all tasks. We include original Graph Transformer and SAN from LRGB (Dwivedi et al., 2022), as well as a variety of recently published models, including MOLTOP baseline (Adamczyk & Czech, 2024), well-tuned classic GNNs (GCN, GatedGCN, GINE) (Tönshoff et al., 2024), and various long-range models such as GRIT (Ma et al., 2023), HDSE (Luo et al., 2024) and S2GCN (Geisler et al., 2024).
Since both datasets are multitask and LightGBM only supports single outputs, we train a separate model per task and combine their predictions. For Peptides-struct, we optimize MAE as a loss function. LRGB results for GNNs are reported for 10 different random seeds. However, LightGBM with default settings is a deterministic classifier, so we don’t report its standard deviation (since it would be zero). In Appendix B, we include results for the Random Forest and Extremely Randomized Trees classifiers, which depend on random seed. However, their standard deviations are very low (<0.1absent0.1<0.1< 0.1 AUPRC).
[…]
ECFP4 fingerprint is roughly equivalent to a 2-layer message-passing GNN. However, it outperforms both local GNNs such as GCN and long-range models like S2GCN and HDSE. We postulate that the reason for this difference is three-fold. First, fingerprints are based on counting discrete substructures, rather than learning continuous graph representations, resulting in a stronger inductive bias. This reduces dependence on training data, as GNNs must learn everything from scratch. Secondly, peptides are built from repeatable fragments, as amino acids are structurally similar and often occur multiple times. Therefore, GNNs may primarily need to recognize that certain small subgraphs, such as the amine group -NH2, frequently occur in peptides. This phenomenon is not indicative of long-range interaction or dependency, but rather reflects the straightforward observation that common subgraphs appear repeatedly across different regions of the input. Third, count-based fingerprints capture the influence of the structure size and frequency of particular substructures. Regression targets in Peptides-struct like inertia mass and length are strongly related to peptide size, and classification labels (e.g. antimicrobial) depend on binding properties of peptides, which also differ with available surface area and thus size. This means that they may not be related to any long-range interactions, but rather local structures and overall peptide size, which can be well approximated by counting subgraphs.
Therefore, our experimental results lead us to believe that Peptides-func and Peptides-struct datasets do not exhibit long-range interactions or require learning long-range dependencies. This has been overlooked in previous works because in (Dwivedi et al., 2022) the authors used only GNNs, and the re-assessment of (Tönshoff et al., 2024) included only quite deep models (6-10 layers) with global topological encodings. This highlights the importance of comparing against established, domain-specific models for fair evaluation. To highlight the impact of using count-based fingerprints, we also compare them with classifiers trained on binary variants, with differences summarized in Table 2. We observe significant performance gains when using count features instead of binary ones. In light of this finding, it is concerning that all the hybrid models in the bioinformatics literature we reviewed rely exclusively on binary fingerprints, especially since count variants incur the same computational cost and result in significantly stronger models.
Table 1: Results on LRGB benchmark.
@@
[…]
@@
We also note the efficiency of our approach. Treating a LightGBM tree node as one parameter, the models for all fingerprints use approximately 300k parameters, in total, well below the budget of 500k set by LRGB. In terms of wall time, computing ECFP fingerprints and training LightGBM takes 19 seconds on a 12-core Intel Core i7-12700KF CPU on Peptides-func. In comparison, training graph transformer or SAN takes up to 60 hours on NVidia A100 GPU (Dwivedi et al., 2022). We provide additional timings in Appendix C.
@@
4.2 Antimicrobial peptides (AMPs) benchmarks
We present results on antimicrobial peptide (AMP) prediction benchmarks. Although the target remains the same (to distinguish peptides with antimicrobial properties), the benchmarks differ in the methodology of selecting positive and negative data, structural diversity, and the way of performing the train-test split (e.g. incorporating sequence homology or clustering). We select 3 benchmarks, covering varied approaches and procedures. Note that many of the datasets used in this section call the data “peptides”, also if those proteins have on average, e.g. 70, 100, or even 200 amino acids. Therefore, to achieve strong performance, fingerprint-based models must be both scalable (owing to their operation at the atomic level) and capable of accurately predicting the properties of larger molecules.
First, we compare our approach with various BERT-like models pretrained for AMP prediction, as evaluated in (Gao et al., 2024). They include AMP peptides from 6 classic datasets, with negative samples taken from UniProt, and additionally filtering the dataset using CD-HIT (Fu et al., 2012) at similarity threshold 40% to remove homology bias. The authors infer only with BERT models, since they have already been pretrained and fine-tuned for AMP prediction. They also note that parts of those datasets may even be in the pretraining datasets, to the advantage of said models on this benchmark. To simulate such pretraining with fingerprints and LightGBM, we use the leave-one-out-dataset strategy. For example, when testing on ADAPTABLE dataset, we merge all other datasets into one training set, and apply CD-HIT-2D (Fu et al., 2012) (threshold 40%) to perform homology filtering, i.e. remove peptides too similar to the ones in the dataset currently used as test set.
[…]
Table 3: Results on AMPs benchmark from (Gao et al., 2024).
@@ 4.2 Antimicrobial peptides (AMPs) benchmarks
Still, those 6 datasets are relatively small. The authors of (Xu et al., 2021) created a unified AMP benchmarking dataset based on 9 commonly used smaller ones (e.g. CAMP and DRAMP), merged and deduplicated using homology-based CD-HIT clustering (Fu et al., 2012). This ensures a high-quality and hard train-test split by avoiding the presence of too similar structures. They benchmarked traditional feature engineering methods, combined with different classifiers (e.g. Random Forest, SVM, kNN).
@@
The comparison is presented in Table 4. Fingerprints achieve SOTA performance, indicating that they perform well at different data sizes. The best previous method, AMPfun (Chung et al., 2019), is a relatively complex pipeline, using binary profiling of amino acid positions, amino acid composition characteristics, and physicochemical descriptors, combined with feature selection. Such approaches can be relatively fragile, as demonstrated by the superior performance of much more straightforward ECFP in both AUROC and MCC metrics.
Table 4: Results on AMPs benchmark from (Xu et al., 2021).
@@
As noted in (Sidorczuk et al., 2022), the negative class sampling has a considerable effect on the model performance. Their AMPBenchmark measures the sensitivity of various AMP models to the choice of data by retraining the models on a large collection of datasets and varying negative class sampling methods. Models vary from sequence-based feature engineering (e.g. AmpGram (Burdukiewicz et al., 2020)), CNN-based deep learning (e.g. Deep-AmPEP30 (Yan et al., 2020)), to multimodal ensembles utilizing physicochemical and amino acid features (e.g. ampir (Fingerhut et al., 2020)).
AMPBenchmark tests the robustness of the models by retraining them on a collection of datasets. The same positive class peptides are combined with 11 different negative class sampling approaches, each repeated 5 times with different data from the UniProt database. This results in 55 datasets, differing in amino acid sequence lengths, sequence similarities, and class balance. Each uses 80-20% train-test split. The results obtained on the AMPBenchmark are summarized in Table 5, including the mean and standard deviation of retraining on different datasets and negative sampling sets. Again, all fingerprint-based models achieve the best results, and with standard deviations considerably lower than other methods. This indicates that feature engineering based on a peptide topological graph may result in more robust and reliable models compared to complex ensembles. Although popular in the bioinformatics literature, they are inherently prone to human bias in the choice of descriptors.
Table 5: Results on AMPs benchmark from (Sidorczuk et al., 2022).
@@
[…]
@@ 4.3 General peptide benchmarks
AutoPeptideML (Fernández-Díaz et al., 2024) consists of 18 datasets. Their positive peptides have been taken from the dataset available in the literature, but the negative samples and the train-test split utilize a novel NegSearch+HP approach to ensure a particularly challenging setup. Negative peptides are taken from APML-Peptipedia, filtered to avoid false negatives, and sampled to have similar sequence length distribution to positive ones. The train-test split uses a homology clustering algorithm to create an out-of-distribution test set, somewhat similar to the scaffold split, popular in molecular property prediction. AutoPeptideML provides results on state-of-the-art PLMs, including LLMs with 3B weights such as Prot-T5-XL (Elnaggar et al., 2022).
The results are summarized in Table 6. Although fingerprint-based approaches do not achieve the highest overall performance, they are close in terms of MCC. Moreover, they are significantly more parameter-efficient, achieving comparable performance with only about 22k parameters, in contrast to PLMs, which can have up to 3 billion weights. In particular, the ECFP-based classifier has better AUROC and MCC than three variants of ESM2 (Lin et al., 2023), which are known to be a particularly powerful protein model, and MCC almost identical to ESM2-650M. The reason for this may be that peptides are very short amino acid sequences, compared to most proteins, which were the bulk of the training and testing data for PLMs. Fingerprints operate on larger molecular graphs, which can also be seen as “higher-resolution" peptide representation than sequences.
Table 6: Results on AutoPeptideML benchmark (Fernández-Díaz et al., 2024).
@@
PeptideReactor (Spänig et al., 2021) is a comprehensive benchmark of 50 datasets, aimed at comparing peptide encoding algorithms. To focus solely on the differences in feature engineering, a Random Forest with 100 trees is consistently used as the classifier. No hyperparameter tuning or class weighting is used for the classifier head, only tuning the feature extraction algorithms is allowed. The authors implemented 48 encodings, based on sequence or structure (using 3D structure descriptors). This benchmark allows us to verify how molecular fingerprints compare to a wide variety of previous peptide encodings from bioinformatics.
@@
For this benchmark, we perform hyperparameter tuning on small grids for fingerprints. We make this exception because, despite its name, many datasets in this benchmark consist of quite large proteins. 17 datasets out of 50 have average sequence lengths of more than 50 amino acids (typical upper limit for peptides), and 8 have average lengths of more than 200. Furthermore, other methods also have their parameters tuned here, in order to fairly assess the capability of the encoding method overall, not its default hyperparameters. With 5-fold CV, we tune: ECFP radius (range [2,4]24[2,4][ 2 , 4 ]), Topological Torsion path length (range [4,6]46[4,6][ 4 , 6 ]), RDKit path length (range [7,9]79[7,9][ 7 , 9 ]). They are still very short-range descriptors, but allowed to capture slightly larger substructures.
@@
Since this is the encoding benchmark, we also check the performance of general “FP encoding“, where we treat the fingerprint type as a hyperparameter. This helps us answer the question "How good are encodings based on the topological graph of a molecule?". This is relevant because the two existing feature groups use amino acid sequence (1D encoding) or spatial peptide structure (3D encoding). Furthermore, this is also more realistic, since we select the best method for each dataset. The results are summarized in Table 7. Due to space limitations, here we report only five best structure- and sequence-based encodings, with detailed results in Appendix D.
[…]
@@
Fingerprints are also highly competitive with top sequence-based methods; notably, the ECFP fingerprint alone secures the 4th position in the benchmark, surpassing 45 other encodings. However, we also noticed quite significant differences between fingerprints depending on the dataset. If we consider the general FP encoding, it achieves the SOTA result on this dataset. This means that fingerprints form robust peptide representations that are efficient across a wide variety of properties and protein sizes. We note that the authors of (Spänig et al., 2021) conclude that structural encodings are inferior to sequence-based ones, both in terms of time required and performance. Fingerprint-based encoding constitutes a counterexample to those findings, albeit based on a topological 2D structure, instead of a spatial one.
Table 7: Results on PeptideReactor benchmark (Spänig et al., 2021).

### v2

@@ 4.1 Long Range Graph Benchmark (LRGB)
To begin, we report the results on the LRGB LRGB benchmark, on both datasets - Peptides-func (classification) and Peptides-struct (regression), in Table 1. Those are multioutput datasets, and we report the metric values averaged from all tasks. We compare to a wide selection of SOTA GNN architectures, including the original Graph Transformer and SAN from LRGB LRGB , as well as a variety of recently published models, e.g. MOLTOP baseline MOLTOP , well-tuned classic GNNs (GCN, GatedGCN, GINE) LRGB_reassessment , and various models with enhanced long-range capabilities such as GRIT GRIT , HDSE HDSE and S2GCN S2GCN .
Since both datasets are multitask and LightGBM only supports single outputs, we train a separate model per task. For Peptides-struct, we optimize MAE as a loss function. LRGB results for GNNs are reported for 10 different random seeds. However, LightGBM with default settings is a deterministic classifier, so we do not report its standard deviation (since it would be zero). We include results for the Random Forest and Extremely Randomized Trees classifiers, which depend on random seed, in the Supplementary Material. Their standard deviations are very low, <0.1<0.1 AUPRC.
The main observation is that all fingerprint-based models achieve state-of-the-art (SOTA) results on both datasets. The strictly local ECFP model, using radius 2 subgraphs, has AUPRC $1.5\%$ higher than the best long-range GNN model, S2GCN. Similarly, Topological Torsion (TT), using only paths of length 4, also outperforms all GNNs. The achievement of SOTA results by these inherently short-range models contradicts the conclusions of LRGB , that long-range dependencies are necessary for peptide function prediction. Instead, our results highlight the critical importance of short-range interactions.
Table 1: The results on LRGB benchmark.
@@
Transformer 63.26 $\pm$ 1.26 0.2529 $\pm$ 0.0016
SAN 64.39 $\pm$ 0.75 0.2545 $\pm$ 0.0012
MOLTOP 64.59 $\pm$ 0.05 -
GraphGPS 65.35 $\pm$ 0.41 0.2500 $\pm$ 0.0005
GINE 66.21 $\pm$ 0.67 0.2473 $\pm$ 0.0017
GatedGCN 67.65 $\pm$ 0.47 0.2477 $\pm$ 0.0009
GCN 68.60 $\pm$ 0.50 0.2460 $\pm$ 0.0007
GraphViT 69.42 $\pm$ 0.75 0.2449 $\pm$ 0.0016
GRIT 69.88 $\pm$ 0.82 0.2460 $\pm$ 0.0012
CRaWl 70.74 $\pm$ 0.32 0.2506 $\pm$ 0.0022
GRED 71.33 $\pm$ 0.11 0.2455 $\pm$ 0.0013
DRew 71.50 $\pm$ 0.44 0.2536 $\pm$ 0.0015
HDSE 71.56 $\pm$ 0.58 0.2457 $\pm$ 0.0013
S2GCN 73.11 $\pm$ 0.66 0.2447 $\pm$ 0.0032
@@
Table 3: The results on BERT-based models benchmark BERT_benchmark .
@@
The ECFP4 fingerprint functions similarly to a two-layer message-passing GNN but outperforms both local models (e.g., GCN) and long-range variants such as S2GCN and HDSE. We attribute this difference to three main factors.
First, fingerprints count discrete substructures rather than learning continuous graph representations, providing a stronger inductive bias and reducing dependence on large training datasets.
Second, peptides consist of repeating fragments—amino acids that are structurally similar and frequently recur. Thus, GNNs mainly need to recognize recurring small subgraphs (e.g., amine groups, -NH2). This phenomenon is not indicative of long-range interaction or dependency, but rather reflects the straightforward observation that common subgraphs appear repeatedly across different regions of the input.
Third, count-based fingerprints naturally encode the influence of structure size and substructure frequency. Regression targets in Peptides-struct like inertia mass and length strongly correlate with peptide size, while in Peptides-func, classification labels (e.g., antimicrobial activity) depend on binding properties linked to surface area and thus size. This means that they may not be related to any long-range interactions, but rather local structures and overall peptide size, which can be well approximated by counting subgraphs.
Our results therefore suggest that the Peptides-func and Peptides-struct datasets do not require modeling long-range interactions. Previous studies may have overlooked this, as LRGB evaluated only GNNs, and LRGB_reassessment used deep models (6–10 layers) with global topological encodings. This highlights the importance of comparing against established, domain-specific models for a fair evaluation. To emphasize the impact of using count-based fingerprints, we also compare them with classifiers trained on binary variants, with differences summarized in Table 2. We observe significant performance gains when using count features instead of binary ones. In light of this finding, it is concerning that all hybrid models in the bioinformatics literature we reviewed rely exclusively on binary fingerprints, despite count variants offering identical computational cost and substantially better performance.
We also note the efficiency of our approach. In terms of wall time, computing ECFP fingerprints and training LightGBM takes 19 seconds on a 12-core Intel Core i7-12700KF CPU on Peptides-func. In comparison, training the graph transformer or SAN takes up to 60 hours on NVidia A100 GPU LRGB . We provide additional timings in the Supplementary Material.
4.2 Antimicrobial peptides (AMPs) benchmarks
We present results on antimicrobial peptide (AMP) prediction benchmarks. Although in all these benchmarks we predict the same property (antimicrobial activity), they differ in their structural diversity, in how positive/negative examples are selected, and in the train–test splitting procedures. We select 3 benchmarks, covering varied approaches and procedures: BERT-based models benchmark BERT_benchmark , XUAMP benchmark Xu_AMP , and AMPBenchmark AMPBenchmark . Note that many of the datasets used in this section contain peptides larger than allowed by typical definition, having on average, e.g. 70, 100, or even 200 amino acids. Therefore, to achieve strong performance, fingerprint-based models must be both scalable (owing to their operation at the atomic level) and capable of accurately predicting the properties of larger molecules.
4.2.1 BERT-based models benchmark
First, we compare our approach with various BERT-based models, pretrained for AMP prediction, as evaluated in BERT_benchmark . They include antimicrobial peptides from 6 classic datasets, with negative samples taken from UniProt, and additionally filter the dataset using CD-HIT CDHIT at the similarity threshold of 40% to remove homology bias. The authors do not perform additional fine-tuning, since BERT models used have already been trained specifically for AMP prediction. They also note that parts of those datasets may even be in the pretraining datasets, to the advantage of said models on this benchmark. To simulate such pretraining with fingerprints and LightGBM we use the leave-one-out-dataset strategy. For example, when testing on ADAPTABLE dataset, we merge all other datasets into one training set, and apply CD-HIT-2D CDHIT (threshold 40%) to perform homology filtering, i.e., remove peptides too similar to the ones in the dataset currently used as test set.
We report the results in Table 3, focusing on the F1 score, since all models obtain very similar AUROC scores, close to 100%. For other metrics, see the Supplementary Material. Fingerprint-based models achieve SOTA results on all datasets, with Topological Torsion (TT) giving the best average F1 score. The difference depends on the dataset, but does not seem to be sensitive to the test dataset size or peptide sizes. For example, ADAPTABLE and DRAMP both have $\approx 4000$ peptides and their average length is $75$ amino acids. However, they differ in difficulty, with fingerprints achieving a $0.3\%$ higher F1 score than the best BERT-based model on ADAPTABLE, but an impressive $12.1\%$ gain on DRAMP. Those results show that the atom-level, short-range fingerprints overall result in excellent models, but the exact data used for training have a considerable impact on the final performance.
4.2.2 XUAMP benchmark
@@
Datasets used in the previous section are popular in AMP research, but are relatively small. The authors of Xu_AMP created a unified AMP benchmarking dataset to combat this problem. Their dataset, also known as XUAMP, is based on 9 commonly used smaller ones (e.g. CAMP and DRAMP), merged and deduplicated using homology-based CD-HIT clustering CDHIT . This prevents data leakage by avoiding the presence of too similar structures in train and test sets, ensuring a realistic evaluation. The authors benchmarked traditional feature engineering methods, combined with different classifiers (e.g. Random Forest, SVM, kNN).
@@
We exactly replicated the original setup and splits of XUAMP, presenting the results in Table 4. Fingerprints achieve SOTA performance, indicating that they perform well at different data sizes. The best previous method, AMPfun AMPfun , is a relatively complex pipeline, using binary profiling of amino acid positions, amino acid composition characteristics, and physicochemical descriptors, combined with feature selection. Such approaches can be relatively fragile, as demonstrated by the superior performance of much more straightforward ECFP in both AUROC and MCC metrics.
Table 4: The results on XUAMP benchmark Xu_AMP .
@@
4.2.3 AMPBenchmark
AMPBenchmark authors AMPBenchmark noted that the negative class sampling has a considerable effect on the model performance. Their benchmark measures the sensitivity of various AMP models to the choice of data by retraining the models on a large collection of datasets and varying negative class sampling methods. Models vary from sequence-based feature engineering (e.g. AmpGram AmpGram ), CNN-based deep learning (e.g. Deep-AmPEP30 DeepAmPEP30 ), to multimodal ensembles utilizing physicochemical and amino acid features (e.g. ampir ampir ).
Robustness of models is tested by retraining them on a collection of datasets. The same positive class peptides are combined with 11 different negative class sampling approaches, each repeated 5 times with different data from the UniProt database. This results in 55 datasets, differing in amino acid sequence lengths, sequence similarities, and class balance. Each uses an 80-20% train-test split. The results obtained on the AMPBenchmark are summarized in Table 5, including the mean and standard deviation of retraining on different datasets and negative sampling sets. Again, all fingerprint-based models achieve the best results, and with standard deviations considerably lower than other methods. This indicates that feature engineering based on a peptide topological graph may result in more robust and reliable models compared to complex ensembles. Although popular in the bioinformatics literature, they are inherently prone to human bias in the choice of descriptors.
Table 5: The results on AMPBenchmark AMPBenchmark .
@@ 4.2.3 AMPBenchmark
AmPEP 61.69 $\pm$ 6.93
iAMP-2L 62.60 $\pm$ 9.51
SVM-LZ 79.71 $\pm$ 2.43
CS-AMPPred 86.69 $\pm$ 4.36
AMAP 89.87 $\pm$ 4.18
Deep-AmPEP30 90.98 $\pm$ 5.61
AmPEPpy 95.35 $\pm$ 2.99
AmpGram 95.56 $\pm$ 2.54
MACREL 96.29 $\pm$ 2.43
MLAMP 96.68 $\pm$ 2.35
ampir 96.71 $\pm$ 2.08
AMPScannerV2 96.79 $\pm$ 2.08
$T$ 97.19 $\pm$ 1.89
RDKit 97.25 $\pm$ 1.84
ECFP 97.37 $\pm$ 1.74
@@ 4.3 General peptide benchmarks
4.3.1 AutoPeptideML
AutoPeptideML AutoPeptideML consists of 18 datasets. Their positive peptides have been taken from datasets available in the literature, while negative peptides are filtered to avoid false negatives and sampled to have similar sequence length distribution to the positive ones. The train-test split uses a homology clustering algorithm to create an out-of-distribution test set, somewhat similar to the scaffold split, popular in molecular property prediction. Those procedures ensure a particularly challenging experimental setup. AutoPeptideML provides results on state-of-the-art PLMs, including LLMs with 3B weights such as Prot-T5-XL ProtTrans .
The results are summarized in Table 6. Although fingerprint-based approaches do not achieve the highest overall performance, they are close both in terms of MCC and AUROC. Moreover, they are significantly more parameter-efficient, achieving comparable performance with only about 22k parameters (counting tree nodes), in contrast to PLMs, which can have up to 3 billion weights. In particular, the ECFP-based classifier has higher AUROC and MCC than three variants of ESM2 ESM2 , which are known to be a particularly powerful protein model, and MCC almost identical to ESM2-650M. The reason for this may be that, compared to most proteins, which were the bulk of the training and testing data for PLMs, peptides are very short amino acid sequences. Fingerprints operate on larger molecular graphs, which can also be seen as “higher-resolution" peptide representation than sequences.
Table 6: The results on AutoPeptideML benchmark AutoPeptideML .
@@ 4.3.2 PeptideReactor
PeptideReactor PeptideReactor is a comprehensive benchmark of 50 datasets, aimed at comparing peptide encoding algorithms. To focus solely on the differences in feature engineering, a Random Forest with 100 trees is consistently used as the classifier. No hyperparameter tuning or class weighting is used for the classifier head, only tuning the feature extraction algorithms is allowed. The authors implemented 48 encodings, based on sequence or structure (using 3D structure descriptors). This benchmark allows us to verify how molecular fingerprints compare to a wide variety of previous peptide encodings from bioinformatics.
@@
For this benchmark, we perform hyperparameter tuning on small grids for fingerprints. We make this exception because, despite its name, many datasets in this benchmark consist of quite large proteins. 17 datasets out of 50 have average sequence lengths of more than 50 amino acids (typical upper limit for peptides), and 8 have average lengths of more than 200. Furthermore, other methods also have their parameters tuned here, in order to fairly assess the capability of the encoding method overall, not its default hyperparameters. With 5-fold CV, we tune: ECFP radius (range $[2,4]$), Topological Torsion path length (range $[4,6]$), RDKit path length (range $[7,9]$). They are still very short-range descriptors, but allowed to capture slightly larger substructures.
@@
Since this is the encoding benchmark, we also check the performance of general “FP encoding“, where we treat the fingerprint type as a hyperparameter. This helps us answer the question "How good are encodings based on the topological graph of a molecule?". This is relevant because the two existing feature groups only use amino acid sequence (1D encoding) or spatial peptide structure (3D encoding). Furthermore, this is also more realistic, since we select the best method for each dataset. The results are summarized in Table 7. Due to space limitations, here we report only five best structure- and sequence-based encodings, with detailed results in the Supplementary Material.
Fingerprints greatly outperform all structural encodings using spatial structure, despite modeling only “flat” molecular graph and short-range interactions. This may mean that those small proteins exhibit quite different biological behavior, compared to large proteins, where the spatial folding structure is known to be very important. Furthermore, molecular graph construction takes just a few seconds, while spatial encodings require a very expensive estimation of the tertiary peptide structure, which takes $\approx 85$ minutes per dataset (see PeptideReactor for details).
@@
Fingerprints are also highly competitive with the top sequence-based methods; notably, the ECFP fingerprint alone secures the 4th position in the benchmark, surpassing 45 other encodings. However, we also noticed quite significant differences between fingerprints depending on the dataset. If we consider the general FP encoding, it achieves the SOTA result on this dataset. This means that fingerprints form robust peptide representations that are efficient across a wide variety of properties and protein sizes. We note that the authors of PeptideReactor conclude that structural encodings are inferior to sequence-based ones, both in terms of time required and performance. Fingerprint-based encoding constitutes a counterexample to those findings, albeit based on a topological 2D structure, instead of a spatial one.
Table 7: The results on PeptideReactor benchmark PeptideReactor .
@@
5 Sequence shuffling analysis
To confirm that short-range dependencies are sufficient for accurate predictions, we perform an additional shuffling experiment. In the training set, we randomly permute (shuffle) a given percentage of amino acids within each sequence, serving as an adversarial perturbation. The test set is kept unchanged. Increasing the shuffling ratio (the percentage of amino acids permuted) progressively disrupts the peptide structure, eliminating its long-range interactions in training data. If the model continues to perform well even on fully randomized sequences, this suggests that peptide properties depend primarily on the molecule substructural composition.
We use the XUAMP, PeptideReactor, and AutoPeptideML benchmarks, reporting the average metric for each using ECFP fingerprints. Results are summarized in Figure 4, with detailed tables provided in the Supplementary Material. The decrease in performance metrics, even for fully randomized sequences, is small—at most about 4%. On XUAMP, the results are nearly identical to the unshuffled baseline. Thus, even when sequences are completely randomized and only very short-range interactions remain, the ECFP fingerprint still achieves strong results. This supports the hypothesis that peptide properties are governed primarily by local, atom-level subgraphs.
Figure 4: Metrics for increasing shuffled amino acids ratio.
6 Data resolution impact
The proposed fingerprint-based methods can be viewed as a higher-resolution extension of a popular baseline, amino acid counts. Instead of counting amino acids or their N-grams at the sequence level, we count substructures directly on the atomic graph. This approach enriches the data representation, particularly in realistic low-data regimes, since the structural encoding inherently incorporates information about amino acids, such as their elemental composition and charge. Consequently, it also captures structural similarities between amino acids, for example, that asparagine and glutamine are more similar to each other than asparagine and proline. In contrast, sequence-level models must learn these relationships implicitly, often requiring more data, as they lack inherent knowledge of amino acid chemistry.
In Table 8, we compare the performance of these two encoding approaches, both using the LightGBM classifier. For amino acid counts, we also include bigram counts, which are commonly employed to improve results. We exclude LRGB due to errors and nonstandard notation in the sequences. Fingerprint-based representations outperform amino acid counts in all cases, by as much as 3.5% on AutoPeptideML and 4.7% on PeptideReactor, highlighting the effectiveness of atom-level data representations for peptide modeling.
Table 8: Average results of amino acid counts and molecular fingerprints per benchmark.
Benchmark Metric Amino acid counts ECFP fingerprint Fingerprint gain
AutoPeptideML AUROC 74.6 78.1 3.5
BERT benchmark AUROC 94.5 95.0 0.5
XUAMP AUROC 74.2 75.0 0.8
AMPBenchmark AUROC 96.9 97.3 0.4
PeptideReactor F1 77.0 81.7 4.7